# Scaling Environments for Organoid Intelligence with LLM-Automated Design and Plasticity-Based Evaluation

### v1

@@ 4.1.2 Environmental scaling
  • Dimensionality: The task can be scaled from a 1D line to a 2D grid or a 3D volume. In a 2D environment, the agent’s state (x, y) could be encoded using separate electrode groups for each axis (e.g., Group C for x-coordinate, Group D for y-coordinate). The action space must expand to four actions (up, down, left, right), requiring a more complex motor decoding scheme, such as using four distinct recording sites or interpreting different spatio-temporal firing patterns from the existing two sites. A pseudocode implementation for this 2D extension is provided in Listing 5.
@@
  • Zone Complexity: The binary safe/aversive zone can be replaced with more complex spatial goals. For instance, the agent could be rewarded for staying on a narrow, winding path or for navigating a simple maze defined by aversive boundaries. This tests the organoid’s ability to learn and represent more intricate spatial relationships. This requires replacing the simple boundary check with a function that evaluates the agent’s position against a predefined geometric structure (see Listing 6).
@@
  • Dynamic Environments: To probe adaptability, the environment can be made non-stationary. The safe zone’s boundary could slowly shift over the course of an experiment, or the exit of a maze could change location after a block of successful trials. This forces the agent to extinguish a previously learned policy and acquire a new one, providing a direct assay for cognitive flexibility (see Listing 7).
@@
Figure 1: Diagram of the Conditional Avoidance environment. The agent, a neural organoid, can move left or right, and will receive negative feedback for entering the aversive zone.
@@
4.2 Environment 2: Goal-seeking in a predator-prey scenario
This task increases complexity by requiring the agent (predator) to actively seek a goal (prey) rather than merely avoiding a negative state. This environment simulates the selective pressures inherent in foraging, a ubiquitous evolutionary challenge that requires an organism to develop efficient strategies for locating and capturing resources (Stephens and Krebs, 1986). The environment can be scaled from a stationary to a moving prey to test for adaptive tracking. A diagram is shown in Figure 2.
4.2.1 Protocol description
  1. 1.
    State Representation and Sensory Encoding: The state is defined by the positions of both the predator and the prey on separate 8-position grids. The agent receives this information via two distinct sensory channels: the prey’s location is encoded by stimulating a specific electrode in group C, while the agent’s own (predator) location is encoded via group D. This separation provides clear exteroceptive (prey location) and proprioceptive (predator location) data streams.
  2. 2.
    Action Decoding and Execution: The motor decoding mechanism is identical to the previous environment, where differential activity between recording groups A and B determines left or right movement for the predator.
  3. 3.
    Feedback and State Transition: Reward is delivered when the predator’s position matches the prey’s position. Upon a successful capture, the prey respawns at a new random location. Punishment, in the form of an unpredictable stimulus, is delivered if the agent fails to capture the prey within a time limit of ZZ timesteps.
4.2.2 Environmental scaling
  • Dimensionality and Search Strategy: Scaling to a 2D or 3D world transforms the task from simple directional movement to a genuine search problem. The agent must learn an efficient strategy (e.g., a patterned or random search) to locate the prey in a larger state space, placing greater demand on the network’s ability to integrate sensory information over time. The sensory and motor schemes must be expanded to accommodate the higher dimensions, as shown in the pseudocode in Listing 8.
  • Introduction of a Predator: A second mobile entity, a predator, can be added to the environment, creating a dual-objective task: approach prey while avoiding the predator. The predator’s location would be encoded on a separate sensory channel (e.g., via a distinct stimulation frequency). Its behavior could range from a random walk to actively tracking the agent. Proximity to the predator would trigger a strong aversive stimulus and could reset the agent’s position (see Listing 9).
  • Population Dynamics: The environment can be populated with multiple prey, forcing the agent to develop a targeting strategy, or multiple predators, requiring more complex avoidance maneuvers.
  • Multi-Organoid Systems: The framework allows for the creation of a simple ecosystem. One organoid could control the prey and another the predator, allowing their behaviors to co-evolve in a competitive dynamic. This can be extended to a linear food chain model (Organoid A is prey to B, B is prey to C) or a cyclical model to study the emergence of hierarchical strategies in a multi-agent biological system. A pseudocode implementation for a two-organoid system is available in Listing  10.
Figure 2: Diagram of the 1D Predator-prey environment. The agent, a neural organoid, can move left or right, and will be rewarded for occupying the same column as the prey.
@@ 4.2.2 Environmental scaling
4.3 Environment 3: Dynamic interception task (Pong)
This experiment replicates the dynamic, real-time control task from (Kagan et al., 2022), providing a benchmark for goal-directed behavior in a more complex, continuous state space. The experimental layout is depicted in Figure 3.
4.3.1 Protocol description
  1. 1.
    Game State Update: The simulation first updates the ball’s position based on its current velocity vector, modeling basic physics within the 2D game world.
  2. 2.
    State Representation and Sensory Encoding: The continuous 2D position of the ball is conveyed to the organoid through a multi-modal encoding scheme. The ball’s horizontal location (X-position) is spatially encoded, where discrete positions maps to specific electrodes across stimulation groups C and D. If each group consists of multiple electrodes, the position will be more precise. Concurrently, the ball’s vertical distance from the paddle (Y-position) is rate-coded, with stimulation frequency being inversely proportional to distance (e.g., 40 Hz when closest, 4 Hz when farthest). This complex sensory input requires the network to integrate two distinct data streams.
  3. 3.
    Action Decoding and Execution: The agent controls its paddle’s vertical movement. Higher aggregate spike rates from recording group A over a 10 ms window move the paddle up, while higher rates from group B move it down.
  4. 4.
    Feedback: A successful interception results in a reward, while a failure triggers an unpredictable punishment.
4.3.2 Environmental scaling
  • Generalization to Other Games: The Pong paradigm can be adapted to other classic arcade games to probe different cognitive functions. For example, Breakout would require precise spatial targeting to aim the ball at specific bricks. Space Invaders would add the complexities of dodging projectiles while timing offensive actions (requiring a third shoot action in the decoded output). A simplified version of Pac-Man would test maze navigation, goal-seeking, and state-dependent behavior (i.e., switching from avoiding ghosts to hunting them). A pseudocode implementation for adapting the task to Breakout is provided in Listing 11.
  • Competitive Multi-Organoid Systems: Pong is an ideal testbed for multi-agent competition. Two separate organoid-MEA systems can be linked to the same game instance, each controlling one of the paddles. This setup allows for the direct study of competitive co-adaptation, where each biological agent must learn and adapt its strategy in real-time response to the actions of the other. The evolution of rally lengths and win/loss rates would serve as powerful metrics for inter-agent learning (see Listing 12).
Figure 3: The Pong environment from (Kagan et al., 2022). The organoid receives complex, multi-modal sensory input about the ball’s position (rate-coded distance, spatially-coded location) and controls the paddle through differential neural activity.
4.4 Automated protocol generation using a large language model
To scale the exploration of experimental parameters and curriculum design, we propose an automated framework where an LLM functions as a meta-controller. The LLM is applied as a tool to orchestrate the agent’s training regimen rather than the agent itself. Manual design of experimental protocols is slow, limited by researcher intuition, and may not efficiently explore the vast parameter space. Automating this process creates a high-throughput system for discovering optimal training strategies and environmental configurations. A pseudocode implementation of this loop is provided in Listing 4.
4.4.1 Framework description
The system operates in a closed loop, iteratively generating, executing, and refining experimental protocols. An empty dataset is created to store all information from each iteration, including generated scripts, execution outputs, performance metrics (e.g., capture rate), neurophysiological responses, and any errors.
  1. 1.
    Prompt Formulation: At the start of an iteration, a detailed prompt is formulated for the LLM via its API. This prompt includes: (1) the high-level experimental objective (e.g., maximize prey capture rate); (2) a description of the available API commands and their valid parameter ranges, covering electrical stimulation (e.g., shape from [bi-phasic, tri-phasic], amplitude from 0.1-20.0 µA, pulse_duration from 50-500 µs, pulse train frequency from 5-100 Hz), microfluidics (flow_rate), and UV uncaging (duration from 100-1000 ms); (3) the current state of the organoid, such as baseline firing rates; and (4) the historical dataset of past iterations, summarizing which protocols succeeded or failed.
  2. 2.
    Protocol Generation: The LLM generates a novel experimental protocol based on the prompt. This can be done in two primary modes: generating a JSON object with specific experimental variables to be plugged into a template script, or generating an entire Python script from scratch that defines the experimental logic. We provide detailed examples of both approaches in Appendix A.3.
  3. 3.
    Validation and Execution: The generated protocol is programmatically validated. A parser checks for syntactic correctness and verifies that all specified parameters fall within safe, predefined ranges. If the protocol is invalid, it is logged to the dataset as erroneous, and the loop returns to the previous step. If valid, the script is executed on the MEA platform. This involves neurosphere stimulation, data collection, administering rewards or punishments, and updating the experiment state.
  4. 4.
    Data Logging: All data from the loop iteration are saved to the dataset. This includes the generated algorithm, the neurosphere’s response, performance metrics, and the resulting experimental state. This comprehensive record is crucial for the refinement stage.
  5. 5.
    Iterative Refinement: Steps 1-4 are repeated for a desired batch size. After each batch, the collected dataset is used to evaluate and refine the LLM’s performance. By identifying patterns in successes and failures, we can improve the LLM through two main avenues:
    • Prompt Engineering: The base prompt is updated. This can involve adding successful examples from the dataset for few-shot prompting, incorporating chain-of-thought reasoning by asking the LLM to explain its choices, or adding explicit constraints based on failed experiments.
    • Fine-tuning: For more significant adaptation, the model can be fine-tuned using the dataset. Successful prompt-protocol pairs serve as training data for supervised fine-tuning, teaching the model to replicate effective strategies. This creates a powerful feedback loop where the LLM becomes an increasingly competent experimental designer over time.
5 Discussion
The framework presented in this paper sits at the intersection of agent-based AI and neuroscience, and its implications warrant careful consideration.
5.1 Broader implications and future work
Treating an organoid as an agent, rather than a passive model, is a paradigm shift. It reframes questions from "What does this neural activity represent?" to "What is this agent trying to do?" This perspective naturally aligns with reinforcement learning principles and predictive coding theories like the Free Energy Principle, suggesting a common computational language for both biological and artificial intelligence.
Looking ahead, the logical progression is to scale the complexity of both the agent and the environment. This includes:
  • Environment Scaling: Moving beyond simple 1D and 2D tasks to more complex, physics-based simulations or even multi-agent scenarios where organoids could interact with each other or with in silico agents.
  • Interface Scaling: Increasing the communication bandwidth through high-density MEAs or integrating optogenetics for precise, cell-type-specific stimulation. This would create a richer observation space and a more expressive action space for the agent.
  • Sim2Real and Embodiment: A major future direction is to bridge the in vitro-to-physical gap. By connecting these systems to simple robotic actuators and sensors, we could create truly embodied biological agents, providing a powerful platform for studying grounded cognition and tackling Sim2Real challenges from a completely new angle.
5.2 Limitations and challenges
Despite its promise, this approach faces significant hurdles. The biological substrate introduces challenges not present in in silico research. Organoids exhibit high variability, have limited lifespans, and their developmental trajectories are difficult to standardize, making reproducibility a key concern. The black box nature of the organoid is profound; while we can measure network-level plasticity, understanding the specific algorithms it implements remains an open question. Furthermore, the information bandwidth of current MEAs is extremely low, potentially constraining the complexity of learnable tasks.
6 Conclusion
This paper presents a framework for the design, scaling, and evaluation of virtual environments for a novel class of biological agent: the neural organoid. By structuring tasks as a curriculum of increasing complexity, from conditional avoidance to dynamic interception, we can systematically probe and induce learning. The proposed methods for sensory encoding, motor decoding, and feedback are grounded in established principles of neuroscience and agent-based learning.
Our key contributions are threefold: (1) we provide a concrete set of scalable environment designs tailored for inducing learning in biological neural networks; (2) we introduce a novel evaluation paradigm for agents, linking behavioral performance to direct measurement of synaptic plasticity; and (3) we propose a forward-looking, LLM-driven methodology for automating and scaling the design of training protocols, which has broad implications for both Organoid Intelligence and AI-driven scientific discovery. This work represents a critical step towards creating more sophisticated and adaptive biological agents, offering a powerful platform to explore the fundamental connections between environment, embodiment, and intelligence.
Appendix A Technical appendices and supplementary material
A.1 Additional pseudocode for initial environments and automation

### v2

@@ 4.1.2 Environmental scaling and long-horizon planning
  • Dimensionality: The task’s complexity can be significantly increased by scaling the world from a 1D line to a 2D grid or a 3D volume. A 2D environment requires a more sophisticated world model representing an (x, y) state, which could be encoded using separate electrode groups or frequency channels for each axis. The action space must expand to four cardinal directions, demanding a more complex motor decoding scheme, such as using four distinct recording sites or interpreting more complex spatio-temporal firing patterns. A pseudocode implementation is provided in Listing 5.
@@
  • Zone Complexity and Planning: The simple binary safe/aversive partition can be replaced with more complex spatial arrangements that require rudimentary planning. For example, the agent could be rewarded for navigating a narrow, winding path or for solving a simple maze defined by aversive boundaries. This challenges the agent to build a world model that represents not just a single boundary, but an extended spatial policy. See Listing 6.
@@
  • Dynamic Environments and Adaptation: To probe the adaptability of the learned world model, the environment can be made non-stationary. The boundary of the safe zone could slowly shift over the course of an experiment, or the goal location in a maze could change after a block of successful trials. This forces the agent to extinguish a previously learned policy and acquire a new one, providing a direct assay for the flexibility and update speed of its internal model (see Listing 7).
@@
Figure 1: Diagram of the Conditional Avoidance environment. The agent, a neural organoid, is embodied in a 1D world. Its decision to move left or right is decoded from neural activity. Entering the aversive zone triggers unpredictable punishment, driving the agent to learn a world model where that region is associated with negative outcomes.
4.2 Environment 2: Goal-seeking in a predator-prey scenario
This task elevates the cognitive demand by requiring the agent (predator) to actively seek a dynamic goal (prey) rather than merely avoiding a static negative state. This environment simulates the fundamental evolutionary pressures of foraging, which necessitates the development of efficient world models for locating and capturing resources under uncertainty (Stephens and Krebs, 1986). A diagram is shown in Figure 2.
4.2.1 Protocol description
  1. 1.
    State Representation and Sensory Encoding: The world state is defined by the positions of both the predator and the prey on separate 8-position grids. The agent receives this information via two distinct sensory channels: the prey’s location (exteroceptive information) is encoded by stimulating a specific electrode in group C, while the agent’s own location (proprioceptive information) is encoded via group D. This separation provides distinct data streams for building a model of itself in relation to the goal.
  2. 2.
    Action Decoding and Execution: The motor decoding mechanism remains identical to the avoidance task, where differential activity between recording groups A and B determines left or right movement for the predator.
  3. 3.
    Feedback and State Transition: A predictable reward is delivered when the predator’s position matches the prey’s position (a successful capture). Upon capture, the prey respawns at a new random location, initiating a new trial. Punishment, in the form of an unpredictable stimulus, is delivered if the agent fails to capture the prey within a time limit of ZZ timesteps, encouraging efficient policies.
4.2.2 Environmental scaling and multi-agent systems
  • Dimensionality and Search Strategy: Scaling the world to 2D or 3D transforms the task from a simple directional problem into a genuine search problem. The agent must learn an efficient search strategy (e.g., patterned, memory-guided, or random) to locate the prey in a vastly larger state space. This places significant demands on the world model’s capacity to integrate sensory information over extended time horizons to infer the target’s location. See Listing 8.
  • Introduction of an Adversary: A second mobile entity, an adversary, can be introduced, creating a multi-objective task: approach prey while avoiding the adversary. The adversary’s location would be encoded on a separate sensory channel (e.g., via a distinct stimulation frequency). Its behavior could range from a random walk to actively tracking the agent, requiring the agent’s world model to account for multiple dynamic entities with different behavioral characteristics. See Listing 9.
  • Multi-Organoid Systems: The framework can be extended to create a simple competitive ecosystem. One organoid could control the prey and a second could control the predator, allowing their policies to co-evolve. This can be further extended to a linear food chain (Organoid A is prey to B, B is prey to C) or a cyclical model, providing a novel platform to study the emergence of complex, hierarchical strategies in a multi-agent biological system. A pseudocode implementation for a two-organoid system is available in Listing 10.
Figure 2: Diagram of the 1D Predator-Prey environment. The agent (predator) must interpret sensory information about its own location and the prey’s location to decide on an action. A successful capture is rewarded, driving the agent to build a world model that supports goal-directed navigation.
4.3 Environment 3: Dynamic interception task (Pong)
This experiment, based on the task from (Kagan et al., 2022), serves as a benchmark for goal-directed behavior in a more complex, continuous-time state space. Success in Pong requires the agent to build a predictive world model of the ball’s trajectory and to generate precisely timed actions to intercept it. The experimental layout is depicted in Figure 3.
4.3.1 Protocol description
  1. 1.
    Game State Update: The simulation first updates the ball’s position based on its current velocity vector, modeling the basic physics of momentum and collision within the 2D game world.
  2. 2.
    State Representation and Sensory Encoding: The continuous 2D position of the ball is conveyed to the organoid through a rich, multi-modal encoding scheme. The ball’s horizontal location (X-position) is spatially encoded, where discrete regions of the X-axis map to specific stimulation electrodes across groups C and D. Concurrently, the ball’s vertical distance from the paddle (Y-position) is rate-coded, with stimulation frequency being inversely proportional to distance (e.g., high frequency when close, low frequency when far). This complex sensory input challenges the agent’s world model to integrate and interpret two distinct, continuous data streams to infer the ball’s state vector.
  3. 3.
    Action Decoding and Execution: The agent controls the vertical movement of its paddle. Higher aggregate spike rates from recording group A over a short time window (e.g., 10 ms) move the paddle up, while higher rates from group B move it down. This provides a continuous-like control mechanism.
  4. 4.
    Feedback: A successful interception (paddle-ball collision) results in a predictable reward, reinforcing the predictive model and the action policy. A failure to intercept (a miss) triggers an unpredictable punishment, signaling a failure in the model’s prediction or the agent’s control.
4.3.2 Environmental scaling and generalization
  • Generalization to Other Games: The core paradigm of Pong can be adapted to other classic arcade games to probe different cognitive functions and world modeling capabilities. For instance, Breakout would require the agent’s world model to incorporate not just the ball and paddle, but also a static array of targets, demanding precise spatial targeting to aim the ball. Space Invaders would introduce the complexities of dodging projectiles while timing offensive actions, requiring the world model to track multiple independent agents. A pseudocode implementation for Breakout is provided in Listing 11.
  • Competitive Multi-Organoid Systems: Pong provides an ideal testbed for multi-agent competition. Two distinct organoid-MEA systems can be linked to the same game instance, each controlling one of the paddles. This setup facilitates the direct study of competitive co-adaptation, where each biological agent must build a world model that includes the policy of its opponent. The evolution of rally lengths, shot patterns, and win/loss rates would serve as powerful metrics for inter-agent learning and strategic model formation (see Listing 12).
Figure 3: Diagram of the Pong environment. The organoid receives complex, multi-modal sensory input about the ball’s position (rate-coded distance, spatially-coded location) and must learn a predictive world model of the ball’s trajectory to control its paddle and achieve a successful interception. The red arrow shows an example trajectory for the orange ball: starting at the top, bouncing off of walls and the agent’s paddle, bouncing off additional walls, and finally passing the agent.
4.4 Generative models for world generation: LLM-driven automated curriculum design
To overcome the limitations of manual experimental design and to systematically explore the vast parameter space of environments and training protocols, we propose an automated framework where an LLM functions as a meta-controller or AI Dungeon Master. This approach directly aligns with the theme of leveraging generative models and LLMs, applying the LLM not as the embodied agent, but as a tool to orchestrate the generation of the agent’s training world and curriculum. This creates a high-throughput system for discovering optimal training strategies and environmental configurations that might be non-obvious to human researchers. A pseudocode implementation of this meta-learning loop is provided in Listing 4.
4.4.1 Framework description
The system operates in a closed loop, iteratively generating, executing, and refining experimental protocols. A central dataset logs all information from each iteration, including the LLM-generated scripts, execution outputs, agent performance metrics (e.g., capture rate, rally length), neurophysiological responses, and any errors encountered.
  1. 1.
    Prompt Formulation: At the start of each iteration, a detailed prompt is programmatically constructed. This prompt provides the LLM with all necessary context, including: (1) the high-level scientific objective (e.g., "Design a protocol to increase the agent’s adaptability to a moving prey"); (2) a comprehensive API specification of available commands and their valid parameter ranges, covering electrical stimulation, microfluidics, and optogenetics; (3) the current physiological state of the organoid (e.g., baseline firing rates, plasticity measurements); and (4) a summary of the historical dataset, providing examples of past protocols that succeeded or failed.
  2. 2.
    Protocol Generation: The LLM processes the prompt and generates a novel experimental protocol. This can be done in two primary modes: generating a structured JSON object with specific experimental variables to be plugged into a template script, or generating an entire Python script from scratch that defines novel experimental logic and curriculum stages. This generative capacity allows the LLM to move beyond simple parameter tuning to design entirely new tasks. Detailed examples are provided in Appendix A.3.
  3. 3.
    Validation and Execution: The generated protocol undergoes automated validation. A parser checks for syntactic correctness and ensures that all specified parameters fall within predefined safety boundaries. Invalid protocols are logged as failures, and the loop continues. Valid protocols are executed on the MEA platform, controlling the closed-loop interaction with the organoid.
  4. 4.
    Data Logging and Analysis: All data from the experimental run, the generated algorithm, the organoid’s neural response, task performance metrics, and the resulting change in physiological state, are logged to the central dataset. This creates a rich, longitudinal record of the agent’s learning journey.
  5. 5.
    Iterative Refinement of the Generative Model: The process repeats. Periodically, the collected dataset is used to refine the LLM’s protocol generation strategy. This can be achieved through:
    • Prompt Engineering: The base prompt is updated with successful examples from the dataset for few-shot learning, or with explicit instructions derived from common failure modes, improving the LLM’s performance in subsequent iterations.
    • Fine-tuning: With a sufficiently large dataset, the model can be fine-tuned on successful prompt-protocol pairs. This process trains the LLM to internalize the features of effective experimental design, creating a powerful feedback loop where the LLM becomes an increasingly sophisticated and specialized scientific discovery tool.
5 Discussion
The framework presented in this paper resides at the confluence of embodied AI, model-based reinforcement learning, and computational neuroscience. By treating a neural organoid as an embodied agent that must learn an internal world model, we open up new avenues for both fundamental scientific inquiry and AI research.
5.1 Broader implications and future work
This work represents a paradigm shift from viewing organoids as passive models of neural circuitry to active agents engaged in goal-directed decision-making. This reframes core questions from "What does this neural activity represent?" to "What predictive model is this agent using to achieve its goals?" This perspective aligns seamlessly with reinforcement learning and predictive coding theories like the Free Energy Principle, suggesting a shared computational language for biological and artificial intelligence. The ability to directly measure the physical implementation of the world model via synaptic plasticity provides a ground truth for learning that is unavailable in purely in silico systems.
The logical progression of this research involves scaling the complexity of both the agent’s embodiment and its environment:
  • Environment Scaling: Moving beyond the simple 2D tasks presented here to more complex, physics-based simulations, partially observable environments, or multi-agent scenarios where organoids can cooperate or compete with each other or with traditional AI agents.
  • Interface Scaling: Increasing the communication bandwidth between the agent and its world through high-density MEAs or integrating optogenetics for cell-type-specific stimulation and recording. A richer I/O channel would enable the agent to form a more detailed world model and execute a more expressive range of actions.
  • Sim-to-Bio-to-Real Embodiment: A crucial future direction is to bridge the in vitro-to-physical gap. By connecting these biological agents to simple robotic actuators (e.g., a wheeled robot) and sensors (e.g., a camera), we could create truly embodied agents that must learn a world model of physical reality. This provides a powerful new platform for studying grounded cognition and tackling sim-to-real challenges from a novel biological perspective, directly addressing aligning simulation with real-world physics.
5.2 Limitations and challenges
Despite its significant promise, this approach is not without its challenges. The biological substrate introduces inherent variability, limited lifespans, and developmental stochasticity, making reproducibility a central concern that must be addressed with robust experimental design and statistical analysis. The organoid itself remains a profound black box; while we can measure network-level plasticity, deciphering the specific computational algorithms it implements to learn and represent its world model is a formidable open problem. Furthermore, the information bandwidth of current MEA technology, while improving, is still orders of magnitude lower than that of biological nervous systems, potentially constraining the complexity of the world models that can be learned.
6 Conclusion
This paper has introduced a comprehensive framework for the design, scaling, and evaluation of virtual environments for a new class of embodied agent: the human neural organoid. By structuring tasks as a coherent curriculum of increasing complexity, from conditional avoidance to goal-seeking and dynamic interception, we can systematically probe and induce the formation of internal world models. Our proposed methods for sensory encoding, motor decoding, and feedback are deeply grounded in the principles of computational neuroscience and model-based reinforcement learning.
Our key contributions are threefold, each directly relevant to the study of embodied world models. First, we provide a concrete set of scalable environment designs tailored to induce and examine learning in biological neural networks. Second, we introduce a novel evaluation paradigm that links an agent’s behavioral performance directly to the physical measurement of synaptic plasticity, offering an unprecedented view into the material basis of a learned world model. Third, we propose a forward-looking, LLM-driven generative methodology for automating and scaling the design of training curricula, which holds broad implications for both Organoid Intelligence and the future of AI-driven scientific discovery. This work represents a critical step toward creating more sophisticated and adaptive biological agents, offering a unique and powerful platform to explore the fundamental connections between environment, embodiment, and intelligence.
Acknowledgments and Disclosure of Funding
We thank FinalSpark for proposing the idea of LLM based automated research.
Appendix A Technical appendices and supplementary material
A.1 Additional pseudocode for initial environments and automation